OpenStudy (anonymous):

i have this question on my homework but its (-5)(-3) = so woud i just subtract or would i add

OpenStudy (tkhunny):

That is a multiplication problem. Why are you just ignoring the parentheses?

OpenStudy (anonymous):

i wasnt

OpenStudy (anonymous):

i just dont know how to solve the question cuz i had one on my test last week and it was (-4) (-4) = and i put 8 and it the answer was 1

OpenStudy (anonymous):

No... The answer was 16.

OpenStudy (anonymous):

How could you have the correct answer as one in your test at school? That's complete rubbish. If you had a line between the brackets then the answer would be 1.

OpenStudy (anonymous):

but on the test it said 16 and 8 were wrong it said 1 was right but it looked like an I

OpenStudy (tkhunny):

I see. It is hard to respond to incorrect information. The ONLY way to get 1 out of (-4)(-4) is by putting DIVISION between them. Maybe it was a misprint, or something. (-4)+(-4) = -8 (-4)-(-4) = 0 (-4)/(-4) = 1 (-4)*(-4) = 16 The ONLY operation assumed by juxtaposition is multiplication. (-5)(-3) = (-5)*(-3) = 15

OpenStudy (anonymous):

Dude... \[(-a)(-b)=-a \times -b\] That's what that means. It's just multiplying numbers. It's the same practicing your "times/muplication tables"

OpenStudy (anonymous):

ok so the correct answer would b 15

OpenStudy (anonymous):

same as*

OpenStudy (anonymous):

-5 time -3 right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

?*

OpenStudy (anonymous):

yess :)

OpenStudy (anonymous):

ok i need help with with simplifying -7[ 6 divided by 2 +3(5-3) and that little 2 at the top outside of the parenthesis

OpenStudy (anonymous):

but using the order of operations

OpenStudy (anonymous):

and this one im completly confused (5+3)3 - 10/2=

OpenStudy (anonymous):

...If you're telling us to use order of operations, why aren't you using it for yourself so you can solve it? Tell me the order of operations from start to finish.

OpenStudy (anonymous):

Paranthesis exponents multiplacation division

OpenStudy (anonymous):

addition and subtraction

OpenStudy (anonymous):

im horrible with math i dnt really know how to solve these so i post them on here so ppl can explain to me how to solve them

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

Mate....This is grade 1 stuff. That's a pretty low standard considering you're in 10th grade!!....

OpenStudy (anonymous):

All I can say is that you got 9 years to catch up and you better hurry up. All I can do is recommend you to get a tutor ASAP. People here on this website can't help you catch up 9 years of maths. You should seriously go and ask your teacher for help. No point going on the internet and asking. That's just useless.

OpenStudy (anonymous):

this is algebra man

OpenStudy (anonymous):

im good at math i just need help on how to do it correctly

OpenStudy (anonymous):

im bad at algebra

OpenStudy (anonymous):

No mate. This is elementary algebra.

OpenStudy (anonymous):

elementary school*

OpenStudy (anonymous):

Too much games and procrastination= this standard.

OpenStudy (anonymous):

there was no algebra in fluttering elementary school

OpenStudy (anonymous):

at least the one i went to

OpenStudy (anonymous):

Dude, you're not doing algebra!! Get it right, you're doing basic arithmetic. Algebra in the most basic definition contains at least one variable or more.

OpenStudy (anonymous):

All you're doing is doing maths that 1st grade students are learning.

OpenStudy (anonymous):

are you gonna just critisize me or r u gonna actually explain how i solve this question

OpenStudy (anonymous):

ik the order of operation i know how to solve it smart retriceim just asking for u to explain it so that ik im doing it right cuz when i do i get stuck on parts

OpenStudy (anonymous):

You haven't giving me an answer yet to your question so I know that you're "doing it right"...

OpenStudy (anonymous):

so I wouldn't know*

OpenStudy (anonymous):

ok so im on -7[ 6 divided by 2 +3(5-3) and that little 2 at the top outside of the parenthesis

OpenStudy (anonymous):

so i i do (5-3) which = 2

OpenStudy (anonymous):

and that little 2 outside the parenthesis i multiply the 2 by 2 which is 4

OpenStudy (anonymous):

this is where im stuck

OpenStudy (anonymous):

cuz the little to is outside the paranthesis

OpenStudy (anonymous):

so would i put 2 inside the paranthesis

OpenStudy (anonymous):

(2) in replace of (5-3)

OpenStudy (anonymous):

Dude, your question is confusing with words. Draw it out with numbers....

OpenStudy (anonymous):